Ask
your own question, for FREE!
Mathematics
12 Online
The odds are 2 to 1 that, when A play B in tennis, A wins. Suppose that A and B play two matches. What is the probability that A wins at least one match?
Still Need Help?
Join the QuestionCove community and study together with friends!
The simple events are E1:{AA} E2:{AB} E3:{BA} E4:{BB}
for example AB means A won the first mach and B wins the second match
My books says that these events are not equally proable, since A wins 2 to 1 eveytime A and B play. They assign the follwing proablilites: P(E1):=4/9, P(E2)=2/9, P(E3)=2/9 and P(E4):=1/9
My question is: How did they get this proablilites?
P(A wins at least 1 match) = 1-P(A loses both matches) =1-1/9 = 8/9
Still Need Help?
Join the QuestionCove community and study together with friends!
okay, but how did they come up with the probablilityes?
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
VERNEISH:
Two objects are moving along separate linear paths where each path is described by position, d, and time, t.
Maci:
what are the best stock apps for people who want to start investing?
Twaylor:
genuienly how do i get better self control i keep watcing youtube and theres this
IzzyCookie:
|dw:1792074056138:dw| Uhm, I don't know where to ask this, but I'm in college welding at my community college, and I need help learning vertical welding for
thehearken:
not sure if this is considered spam, im sorry if it is, but does anyone have tips
1 day ago
1 Reply
0 Medals
6 days ago
3 Replies
2 Medals
1 week ago
3 Replies
0 Medals
3 days ago
9 Replies
1 Medal
1 week ago
2 Replies
1 Medal